OpenStudy (anonymous):

What is the slope of the line through the points (2, 0) and (3, -3)? m = _?_

OpenStudy (anonymous):

m= y1-y2 --------- x1-x2

OpenStudy (anonymous):

y1-y2 is the value of the slope?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

(-3-0)/(3-2)= -3/1= -3

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

THANKS!!!